# Setting Limits on Blazar-Boosted Dark Matter with Xenon-based Detectors

### v1

@@ II.3 Dark Matter Density profile
where $\alpha=(9-2\gamma)/(4-\gamma)$ and $R_{\text{spike}}$ is the radial extension of the spike. For $\gamma=1$ as in the NFW profile, the GS spike has slope $\alpha_{{\mathrm{GS}}}=7/3$. The existence of spikes around super massive BHs has recently received observational support from Ref. [42], which found evidence for a GS spike at low radii and NFW behavior at large radii for the supermassive BH in blazar OJ 287.
However, numerical simulations generally predict that the spike is modified when additional physical processes are included, such as baryonic effects, non-circular orbits, and displacement of the BH from the halo center [43]. Stellar scattering effects on the dark matter density spike have been extensively studied using the Fokker-Planck equation [38], two-fluid approximations, and N-body simulations [40, 39].
@@
where $G$ is Newton’s gravitational constant. For the stellar velocity dispersion we have the following relation in terms of $M_{{\mathrm{BH}}}$ [44]:
@@
which implies that for our value of $M_{{\mathrm{BH}}}$ (Tab. 1) $R_{\text{infl}}\simeq 49\,\text{pc}$.
@@
where $m_{\chi}$ is the mass of the DM particle and $\langle\sigma_{\chi\chi}v_{\text{rel}}\rangle$ is the thermally averaged annihilation cross section times relative velocity. For this analysis, we assume self-annihilation effects are negligible within the radius of influence, effectively taking $\langle\sigma_{\chi\chi}v_{\text{rel}}\rangle\to 0$. This assumption maximizes the dark matter density available for boosting but should be reconsidered for models with significant self-interaction cross-sections.
3-Zone Model. The DM profile used in this work, $\rho_{\text{3ZM}}(r)$, implements a 3-zone profile similar to that used in Ref. [45]. It includes baryonic effects up to $R_{\text{infl}}$, a GS spike up to $R_{\text{spike}}$, and transitions to NFW at larger radii. The full profile is given as:
@@ III.1 BBDM flux
To calculate the flux of blazar-boosted dark matter reaching Earth, we need to model the scattering process between DM particles and protons in the blazar jet. The same approach applies to DM-nucleus interactions at the detector. This scattering depends on the internal structure of the target particles, which is captured by form factors that provide approximations of their charge distributions.
For DM-proton scattering, we use the dipole form factor
@@
where $Q^{2}$ is the squared momentum transfer and $\Lambda\simeq 0.84\,\text{GeV}$ for the proton. While ideally we would use a form factor derived from the weak charge distribution, accurate measurements of this would require precise neutrino-proton scattering data, which is not currently available.
For interactions with heavy nuclei in xenon detectors, we use the standard Helm form factor, which accounts for the spatial distribution of nucleons within the nucleus and is the established choice for xenon-based dark matter experiments [46, 47].
If the cross section $\sigma_{\chi p}$ of the elastic scattering between the DM particle $\chi$ and the proton is assumed to be isotropic in the center-of-mass (c.m.) rest frame and independent of the c.m. energy, the BBDM flux $\Phi_{\chi}$ for a dark matter particle with mass $m_{\chi}$ is given by [7, 8]:
@@
where $\Sigma_{{\mathrm{DM}}}^{\text{tot}}=\Sigma_{{\mathrm{DM}}}(r_{\mathrm{max}})$ is the total dark matter column density along the line of sight through the blazar jet, $r_{\mathrm{max}}$ is the maximal radial extension of the blazar jet, $d_{L}$ is the luminosity distance, $\phi_{s}$ is the azimuthal angle of scattering with respect to the l.o.s., $T_{p}^{\mathrm{min}}(T_{\chi})$ is the minimal kinetic energy required to have an outgoing DM particle with kinetic energy $T_{\chi}$, $T_{p}^{\mathrm{max}}$ is the maximal energy of the protons available in the blazar jet, and $T_{\chi}^{\mathrm{max}}(T_{p})$ is the maximal kinetic energy that the DM has after a forward scattering with a proton of energy $T_{p}$. The proton spectrum needs to be evaluated at the proper angle that leads to scattering of DM towards the Earth. Consequently, since, in our case, $\theta_{\text{l.o.s.}}\neq 0$, the proton spectrum inherits a non-trivial dependence on $\phi_{s}$ and $T_{p}$ (see, e.g., Ref. [8] for more details). Using the attenuation and detector response described in the following sections, we found that the signal in an lXe TPC is not affected if we cut off the BBDM spectrum from the blazar above $\sim$1\text{\,}\mathrm{GeV}$\$, meaning that the results will not be affected by propagator effects at or above the weak scale.
@@ III.2 Earth attenuation
We model this attenuation using the analytical energy-loss method from Eq.11 of Ref.[48]. While more detailed Monte Carlo simulations using Darkprop [49] are available, we use the analytical approach due to computational constraints. However, the choice of attenuation method is not the leading source of uncertainty in our limits; the dark matter density profile has a much greater impact on the final constraints.
Figure 2: Comparison of nuclear recoil energy spectra in liquid xenon for a $46.4\text{\,}\mathrm{keV}\text{\,}{\mathrm{\text{$c$}}}^{-2}$ WIMP boosted by TXS 0506+056. In black is the expected energy deposition spectrum in liquid xenon, while in blue is the signal from the LZ EFT analysis that best fits the blazar-boosted energy deposition, a $1\text{\,}\mathrm{TeV}$ WIMP coupling to nuclear spin ($\mathcal{O}_{4}$). Green lines show the efficiency for the XENON1T WIMP search reaching $\sim$50 $\text{\,}\mathrm{keV} [50$], and for the LZ EFT analysis where the energy range is extended to above $250\text{\,}\mathrm{keV}$ [51]. The comparison illustrates how blazar-boosted dark matter can produce signals in energy ranges that extend beyond typical WIMP search windows.
@@ V Statistical interpretation using public data
Data from liquid xenon time-projection-chambers (LXe TPCs) are interpreted in terms of specific dark matter-xenon interaction models, most commonly assuming a spin-independent WIMP-nucleon interaction, with a standard halo model velocity distribution [47] and considering nuclear recoils between $3\text{\,}\mathrm{keV}$ to $\sim$40\text{\,}\mathrm{keV}$\$ as shown in the Fig. 2 line for XENON1T.
The simplest method for computing approximate limits is what we refer to as “rate-matching,” following the approach in Ref. [7]. This method finds the cross-section required for the blazar-boosted signal rate in the detector to match the published WIMP limit event rate. This approach takes almost no spectral information into account other than detector efficiency. In this paper, we improve accuracy and sensitivity by incorporating as much spectral information as possible into the search.
XENON1T case study. The XENON1T collaboration has provided a complete statistical model of their tonne-year search  [50, 52]. We use this model to compute constraints on BBDM (Fig. 3) and to validate the “rate-matching” approach. We find qualitative agreement between the two methods, with deviations of around a factor of 2 in the rightmost corner of the exclusion region. These larger deviations occur where attenuation most strongly affects the recoil spectra.
LZ case study. BBDM may yield significantly higher recoil energies than the typical WIMP signature. The LZ collaboration has developed a nuclear recoil model to extend their search region to almost $300\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ [51] to constrain effective field theory (EFT) interactions. Combined with LZ’s generally higher sensitivity to WIMPs, this extended energy range significantly improves constraints on boosted DM, such as in Ref. [6]. To incorporate spectral information from LZ, we use $\chi^{2}$ minimization to find the EFT operator from Ref. [51] that best matches each Earth-attenuated BBDM recoil spectrum. We then find the BBDM cross-section that produces the rate excluded by LZ. We perform this matching and compute exclusions across a grid of BBDM masses and cross-sections, with results shown in Fig. 3. We validated this matching approach by comparing it with a more conservative approach where instead of the best-match scale we demand that the BBDM spectrum at the limit exceeds the matched spectrum in a recoil energy region that contains $90\%$ of the signal.
@@
Figure 3: Constraints on the DM cross-section as function of WIMP mass for blazar-boosted dark matter. The Blue line shows the constraint using the XENON1T open likelihood [52, 50], while the red lines indicate recasted limits from the LZ EFT search [51]. Full lines illustrate the constraint using the three-zone model DM column density, while the dashed and dotted lines show the constraints for the NFW and GS DM density model, respectively. The upper edge of the constrained region is due to the attenuation of the DM flux by the Earth.
@@ VI Conclusions
Our key finding is that astrophysical uncertainties, particularly the dark matter density profile near the supermassive black hole, represent the primary limitation on sensitivity rather than detector systematics. The constraints vary by more than an order of magnitude depending on the adopted density profile, highlighting the critical importance of inner halo structure modeling. We demonstrate how to recast direct detection results using full spectral information rather than simple rate-matching, and find that extending the nuclear recoil energy range, as demonstrated by LZ’s coverage up to $\sim$$300\text{\,}\mathrm{keV}$ compared to XENON1T’s $\sim$$50\text{\,}\mathrm{keV}$ cutoff, provides substantial improvements in sensitivity since boosted dark matter deposits significantly higher energies than typical halo WIMPs.
These results establish blazar-boosted dark matter as a promising complementary probe for sub-GeV dark matter, extending sensitivity to mass ranges inaccessible to conventional halo searches. However, blazar-boosted searches are fundamentally limited by astrophysical uncertainties for now, whereas conventional galactic halo searches are primarily constrained by detector sensitivity. Our work motivates both improved astrophysical modeling of galactic centers and extension to multiple blazar sources. We encourage direct detection experiments to publish their full results in formats that facilitate recasting for alternative dark matter scenarios. The framework developed here can be readily applied to next-generation detectors such as DARWIN, and the strong coupling between dark matter searches and galactic structure studies opens new avenues for multi-messenger constraints on both dark matter physics and the central regions of massive galaxies.
Acknowledgements

### v3

@@ II.3.2 Gondolo-Silk
where $\alpha=(9-2\gamma)/(4-\gamma)$ and $R_{\text{spike}}$ is the radial extension of the spike. For $\gamma=1$ as in the NFW profile, the GS spike has slope $\alpha_{{\mathrm{GS}}}=7/3$. The existence of spikes around super massive BHs has recently received observational support from Ref. [chan2024first], which found evidence for a GS spike at low radii and NFW behavior at large radii for the supermassive BH in blazar OJ 287.
However, numerical simulations generally predict that the spike is modified when additional physical processes are included, such as baryonic effects, noncircular orbits, and displacement of the BH from the halo center [Ullio2001]. Stellar scattering effects on the dark matter density spike have been extensively studied using the Fokker-Planck equation [Gnedin2004Aug], two-fluid approximations, and $N$-body simulations [Shapiro2022Aug, Merritt2007Feb].
@@
where $G$ is Newton’s gravitational constant. For the stellar velocity dispersion we have the following relation in terms of $M_{{\mathrm{BH}}}$ [Kormendy:2013dxa]:
@@
which implies that for our value of $M_{{\mathrm{BH}}}$ (Table 1) $R_{\text{infl}}\simeq 49\,\text{pc}$.
@@
where $m_{\chi}$ is the mass of the DM particle and $\langle\sigma_{\chi\chi}v_{\text{rel}}\rangle$ is the thermally averaged annihilation cross section times relative velocity. For this analysis, we assume self-annihilation effects are negligible within the radius of influence, effectively taking $\langle\sigma_{\chi\chi}v_{\text{rel}}\rangle\to 0$. This assumption maximizes the dark matter density available for boosting but should be reconsidered for models with significant self-interaction cross sections.
II.3.3 3-Zone Model
The DM profile used in this work, $\rho_{\text{3ZM}}(r)$, implements a 3-zone profile similar to that used in Ref. [Cline:2023tkp]. It includes baryonic effects up to $R_{\text{infl}}$, a GS spike up to $R_{\text{spike}}$, and transitions to NFW at larger radii. The full profile is given as:
@@ III.1 BBDM flux
To calculate the flux of blazar-boosted dark matter reaching Earth, we need to model the scattering process between DM particles and protons in the blazar jet. The same approach applies to DM-nucleus interactions at the detector. Throughout this work, for baseline demonstration, we follow the result-reporting conventions of dark matter direct detection experiments [Baxter:2021pqo], modeling the DM-proton scattering with a fixed total cross section that is energy-independent. For DM-proton scattering, we use the dipole form factor
@@
where $Q^{2}$ is the squared momentum transfer and $\Lambda\simeq 0.84\,\text{GeV}$ for the proton. While ideally we would use a form factor derived from the weak charge distribution, accurate measurements of this would require precise neutrino-proton scattering data, which is not currently available. In our kinematic regime the momentum transfer is much smaller than $\Lambda$, so $G_{D}(Q^{2})\approx 1$ and the dipole form factor introduces no relevant energy dependence.
For interactions with heavy nuclei in xenon detectors, we use the standard Helm form factor, which accounts for the spatial distribution of nucleons within the nucleus and is the established choice for xenon-based dark matter experiments [Lewin:1995rx, Baxter:2021pqo].
If the cross section $\sigma_{\chi p}$ of the elastic scattering between the DM particle $\chi$ and the proton is assumed to be isotropic in the center-of-mass (c.m.) rest frame and independent of the c.m. energy, the BBDM flux $\Phi_{\chi}$ for a dark matter particle with mass $m_{\chi}$ is given by [Wang:2021jic, Granelli:2022ysi]:
@@
where $\Sigma_{{\mathrm{DM}}}^{\text{tot}}=\Sigma_{{\mathrm{DM}}}(r_{\mathrm{max}})$ is the total dark matter column density along the line of sight through the blazar jet, $r_{\mathrm{max}}$ is the maximal radial extension of the blazar jet, $d_{L}$ is the luminosity distance, $\phi_{s}$ is the azimuthal angle of scattering with respect to the l.o.s., $T_{p}^{\mathrm{min}}(T_{\chi})$ is the minimal kinetic energy required to have an outgoing DM particle with kinetic energy $T_{\chi}$, $T_{p}^{\mathrm{max}}$ is the maximal energy of the protons available in the blazar jet, and $T_{\chi}^{\mathrm{max}}(T_{p})$ is the maximal kinetic energy that the DM has after a forward scattering with a proton of energy $T_{p}$. The proton spectrum needs to be evaluated at the proper angle that leads to scattering of DM toward the Earth. Consequently, since, in our case, $\theta_{\text{l.o.s.}}\neq 0$, the proton spectrum inherits a nontrivial dependence on $\phi_{s}$ and $T_{p}$ (see, e.g., Ref. [Granelli:2022ysi] for more details).
Using the attenuation and detector response described in the following sections, we examined how the high energy part of the BBDM spectrum contributes to the predicted BBDM rate at the detector. We compared the recoil spectrum obtained by integrating the flux to arbitrarily high energies with the spectrum obtained when truncating it at $T_{\chi}\simeq 1$ GeV. Since the difference is negligible, we truncated the BBDM flux to include energies only up to $T_{\chi}=1$ GeV. This can be understood intuitively: because the DM mass is many orders of magnitude smaller than the xenon nuclear mass, even relativistic DM can transfer only a small fraction of its kinetic energy in an elastic collision. As a result, the maximum recoil energy saturates at a few hundred keV, and the very high energy part of the BBDM flux does not produce additional detectable events in liquid xenon. This conclusion is driven entirely by recoil kinematics and the steeply falling BBDM flux, and therefore remains valid even if the DM–nucleus cross section has a mild energy dependence.
@@ III.2 Earth attenuation
We model this attenuation using the analytical energy-loss method from Eq. (11) of Ref.[Bringmann:2018cvk]. While more detailed Monte Carlo simulations using Darkprop [Xia:2021vbz] are available, we use the analytical approach due to computational constraints. However, the choice of attenuation method is not the leading source of uncertainty in our limits; the dark matter density profile has a much greater impact on the final constraints.
Figure 2: Comparison of nuclear recoil energy spectra in liquid xenon for a $46.4\text{\,}\mathrm{keV}\text{\,}{\mathrm{\text{$c$}}}^{-2}$ WIMP boosted by TXS 0506+056. In black is the expected energy deposition spectrum in liquid xenon, while in blue is the signal from the LZ EFT analysis that best fits the blazar-boosted energy deposition, a $1\text{\,}\mathrm{TeV}$ WIMP coupling to nuclear spin ($\mathcal{O}_{4}$). Green lines show the efficiency for the XENON1T WIMP search reaching $\sim$$50\text{\,}\mathrm{keV}$ [XENON:2018voc], and for the LZ EFT analysis where the energy range is extended to above $250\text{\,}\mathrm{keV}$ [LZ:2023lvz]. The comparison illustrates how blazar-boosted dark matter can produce signals in energy ranges that extend beyond typical WIMP search windows.
@@ V Statistical interpretation using public data
Data from liquid xenon time-projection-chambers (LXe TPCs) are interpreted in terms of specific dark matter-xenon interaction models, most commonly assuming a spin-independent WIMP-nucleon interaction, with a standard halo model velocity distribution [Baxter:2021pqo] and considering nuclear recoils between $3\text{\,}\mathrm{keV}$ to $\sim$40\text{\,}\mathrm{keV}$\$ as shown in the Fig. 2 (see the line for XENON1T).
The simplest method for computing approximate limits is what we refer to as “rate-matching,” following the approach in Ref. [Wang:2021jic]. This method finds the cross section required for the blazar-boosted signal rate in the detector to match the published WIMP limit event rate. This approach takes almost no spectral information into account other than detector efficiency. In this paper, we improve accuracy and sensitivity by incorporating as much spectral information as possible into the search.
We use XENON1T and LZ public data to derive constraints on the DM–nucleon cross section.
XENON1T case study. The XENON1T collaboration has provided a complete statistical model of their tonne-year search  [XENON:2018voc, xenon1tbinference]. We use this model to compute constraints on BBDM (Fig. 3) and to validate the “rate-matching” approach. We find qualitative agreement between the two methods, with deviations of around a factor of 2 in the rightmost corner of the exclusion region. These larger deviations occur where attenuation most strongly affects the recoil spectra.
LZ case study. BBDM may yield significantly higher recoil energies than the typical WIMP signature. The LZ collaboration has developed a nuclear recoil model to extend their search region to almost $300\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ [LZ:2023lvz] to constrain effective field theory (EFT) interactions. Combined with LZ’s generally higher sensitivity to WIMPs, this extended energy range significantly improves constraints on boosted DM, such as in Ref. [LZ:2025iaw]. To incorporate spectral information from LZ, we use $\chi^{2}$ minimization to find the effective field theory (EFT) operator from Ref. [LZ:2023lvz] that best matches each Earth-attenuated BBDM recoil spectrum. We then find the BBDM cross section that produces the rate excluded by LZ. We perform this matching and compute exclusions across a grid of BBDM masses for the assumed BBDM model and cross sections, with results shown in Fig. 3. We validated this matching approach by comparing it with a more conservative approach where instead of the best-match scale we demand that the BBDM spectrum at the limit exceeds the matched spectrum in a recoil energy region that contains $90\%$ of the signal.
@@
About the constant cross section. In setting exclusion limits we adopted a constant–cross section approximation (as in Ref. [Bringmann:2018cvk]). In reality, the full relativistic DM–nucleon scattering cross section retains energy dependence [Dent:2019krz, de2025boosted]—even in the heavy-mediator regime. This behavior is explicitly shown in Refs. [ema2021neutrino, bell2024cosmic], where mediator masses of order a GeV still lead to clear departures from an energy–independent interaction.
[…]
At small cross sections—where Earth attenuation is negligible—the constant cross–section approximation is conservative: any realistic energy dependence would increase the predicted scattering rate and therefore strengthen (not weaken) the limits. In practice, these theoretical uncertainties are much smaller than the astrophysical uncertainties in the dark–matter density profile near the black hole, which dominate the variation in the final constraints.
Figure 3: Constraints on the DM cross section as function of WIMP mass for blazar-boosted dark matter. The blue line shows the constraint using the XENON1T open likelihood [xenon1tbinference, XENON:2018voc], while the red lines indicate recasted limits from the LZ EFT search [LZ:2023lvz]. Full lines illustrate the constraint using the three-zone model DM column density, while the dashed and dotted lines show the constraints for the NFW and GS DM density model, respectively. The upper edge of the constrained region is due to the attenuation of the DM flux by the Earth. We show the same limits for different dark-matter density profiles to highlight that the choice of halo model can strongly affect the resulting constraints.
@@ VI Conclusions
Our key finding is that astrophysical uncertainties, particularly the dark-matter density profile near the supermassive black hole, represent the primary limitation on sensitivity rather than detector systematics. The constraints vary by more than an order of magnitude depending on the adopted density profile, highlighting the critical importance of inner halo structure modeling. We demonstrate how to recast direct detection results using full spectral information rather than simple rate-matching, and find that extending the nuclear recoil energy range, as demonstrated by LZ’s coverage up to $\sim$$300\text{\,}\mathrm{keV}$ compared to XENON1T’s $\sim$$50\text{\,}\mathrm{keV}$ cutoff, provides substantial improvements in sensitivity since boosted dark matter deposits significantly higher energies than typical halo WIMPs. We also compared the full–detector–response method with a simple rate-matching approach and found that the two generally agree at the qualitative level, differing noticeably only in regions where attenuation becomes important.
These results establish blazar-boosted dark matter as a complementary probe for sub-GeV dark matter, extending sensitivity to mass ranges inaccessible to conventional halo searches. However, blazar-boosted searches are fundamentally limited by astrophysical uncertainties for now, whereas conventional galactic halo searches are primarily constrained by detector sensitivity. Our work motivates both improved astrophysical modeling of galactic centers and extension to multiple blazar sources. We encourage direct detection experiments to publish their full results in formats that facilitate recasting for alternative dark matter scenarios. The framework developed here can be readily applied to next-generation detectors such as DARWIN, and the strong coupling between dark matter searches and galactic structure studies opens new avenues for multimessenger constraints on both dark matter physics and the central regions of massive galaxies.
Acknowledgments
@@
Data availability
The data that support the findings of this article are openly available [xenon1tbinference, LZ:2023lvz].